OpenStudy (anonymous):

If both the sequence (nan) and the series n(an - an+1) converge, prove that the series (an) also converges. hint: consider the sequence of partial sums and use the ALT for sequences.

OpenStudy (anonymous):

dizzy :D

OpenStudy (anonymous):

The sequence \(\large na_n\) converges and the series \[\Large \sum_{n=k}^\infty n(a_n-a_{n+1})\] also converges... so what are we trying to prove again? ^.^

OpenStudy (anonymous):

that (an) converges as well

OpenStudy (anonymous):

as a sequence or as a series?

OpenStudy (anonymous):

oo series sorry i left that part out of the ques

OpenStudy (anonymous):

You want to prove that \[\Large \sum_{n=k}^\infty a_n\] is also convergent?

OpenStudy (anonymous):

says with n=1

OpenStudy (anonymous):

yes*

OpenStudy (anonymous):

looks daunting. but anyway, let's consider the sequence of partial sums ^.^ \[\Large s_n:=\Large \sum_{k=1}^nk(a_n-a_{n+1})\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whoops, i meant \[\Large s_n:=\Large \sum_{k=1}^nk(a_k-a_{k+1})\]

OpenStudy (anonymous):

np i understood what u meant

OpenStudy (anonymous):

So, can you derive a nicer expression for \(\large s_n\)? One that doesn't involve that big sigma...

OpenStudy (anonymous):

i dont follow?

OpenStudy (anonymous):

Here, I'll show you... what's \(\large s_1\) equal to?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

or 0

OpenStudy (anonymous):

no no... Just substitute, n = 1 \[\Large s_1 = \sum_{k=1}^{1}k(a_k-a_{k+1})\]

OpenStudy (anonymous):

or wait

OpenStudy (anonymous):

I believe this is equal to \[\Large a_1 - a_2\]?

OpenStudy (anonymous):

is it a1 - a2?

OpenStudy (anonymous):

ok yes sorry

OpenStudy (anonymous):

Okay, kick it up a notch ^.^ What about \(\large s_2\)?

OpenStudy (anonymous):

2(a2-a3)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No, that's just the second term of the sequence, but if you consider it a series, you have to add it to the previous term... \[\Large a_1 - a_2 + 2(a_2-a_3) \]

OpenStudy (anonymous):

oo ok so then s3= a1-a2+2(a2-a3)+3(a3-a4)?

OpenStudy (anonymous):

I think we can simplify a2 before we proceed to a3\[\Large a_1 - a_2 + 2(a_2-a_3) = a_1+a_2 - 2a_3\]

OpenStudy (anonymous):

ok makes sense

OpenStudy (anonymous):

Now you can add the third term. s3 = ?

OpenStudy (anonymous):

a1+a2+a3-3a4?

OpenStudy (anonymous):

That's right LOL

OpenStudy (anonymous):

Can you now find a nice representation for sn?

OpenStudy (anonymous):

sn=a1+a2+a3....+an

OpenStudy (anonymous):

or ..+an+1?

OpenStudy (anonymous):

<whistles>

OpenStudy (anonymous):

Allow me to make a list of what we know ^_^ \[\Large s_1 = a_1 - a_2\]\[\Large s_2 = a_1+a_2 - 2a_3\]\[\Large s_3 = a_1+a_2+a_3-3a_4\]\[\Large s_4 = a_1+a_2+a_3+a_4-4a_5\] and so on, right?

OpenStudy (anonymous):

oo right

OpenStudy (anonymous):

im drawing a blank on how to put it into one summation

OpenStudy (anonymous):

Well, how about this...\[\Large s_1 = a_1+a_2 -2a_2\]\[\Large s_2= a_1+a_2+a_3 - 3a_3\]\[\Large s_3 = a_1+a_2+a_3 + a_4 -4a_4\]\[\Large s_4 = a_1+a_2+a_3+a_4 +a_4 - 5a_5\] A little modification.

OpenStudy (anonymous):

sorry, typo on the last part \[\Large s_4 = a_1+a_2+a_3+a_4 +a_\color{red}5 - 5a_5\]

OpenStudy (anonymous):

ok i was just gunna ask that lol

OpenStudy (anonymous):

As for compressing them into one summation, you're severely lacking imagination XD

OpenStudy (anonymous):

\[\Large s_{n-1} = \left[\sum_{k=1}^{n}a_k\right]-na_n\]

OpenStudy (anonymous):

ooo ok got it

OpenStudy (anonymous):

But you know sn is convergent right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Because it's the series of partial sums

OpenStudy (anonymous):

and so is nan

OpenStudy (anonymous):

We can fix this so that it becomes \[\Large s_{n-1}+na_n=\sum_{k=1}^na_k \]

OpenStudy (anonymous):

That was just algebra... are you still with me? :)

OpenStudy (anonymous):

yes sorry im just looking over the steps as we are going

OpenStudy (anonymous):

oh okay take your time ^.^

OpenStudy (anonymous):

ok go it

OpenStudy (anonymous):

got*

OpenStudy (anonymous):

This part says it all: \[\Large s_{n-1}+na_n=\sum_{k=1}^na_k\] You know THIS converges (this is simply the sequence of partial sums of a known convergent series) \[\Large \color{green}{s_{n-1}}+na_n=\sum_{k=1}^na_k\], so suppose it converges to some value s You also know that this converges (it was given in your question) suppose it converges to some value x.\[\Large s_{n-1}+\color{red}{na_n}=\sum_{k=1}^na_k\]

OpenStudy (anonymous):

So if we take the limit on both sides as n goes to infinity... \[\Large \lim_{n\rightarrow \infty}\left (s_{n-1}+na_n\right)=\lim_{n\rightarrow\infty }\sum_{k=1}^na_k\] We get \[\Large \color{green}s+\color{red}x=\color{blue}{\sum_{k=1}^\infty a_k}\]

OpenStudy (anonymous):

I wouldn't mind a little applause ^.^

OpenStudy (anonymous):

lol ur a life saver!..i never understand this stuff but when u take me thru it step by step i get it more..thank you soo much!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i have another question if u have time lol

OpenStudy (anonymous):

This was a mouthful. go ahead :D

OpenStudy (anonymous):

this one is even harder lol..ok il close this one and post another